Jane's mother is 42 years older than her. The square of Janes age is equal to her mothers age. How old is Jane andher mother?
Answer by ankor@dixie-net.com(22740)

Jane's mother is 42 years older than her.
m = j + 42
The square of Jane's age is equal to her mothers age.
m = j^2
Therefore
j^2 = j + 42
j ^2 - j - 42 = 0; a quadratic equation
factors to
(j-7)(j+6) = 0
The positive solution (Jane has to be positive)
j = 7 years old
How old is Jane and her mother?
7 and 49